OpenStudy (anonymous):

Create two trinomials that you can add together to get 4x - 8. Explain how it is possible to combine (add) two trinomials and end up with a binomial answer. WILL GIVE MEDAL!!! BUT I NEED IT EXPLAINED

OpenStudy (anonymous):

let the two trinomials be: trinomial1: \(\ ax^2+bx+c \) and trinomial2: \(\ dx^2+ex+f \) ok so far?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

i have to turn this is sooo thats all im supposed to turn in?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

no... im explaining how to get the answer....

OpenStudy (anonymous):

can you just type it all out then i write it all down? thats muchhhh easierrr @dpaInc

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

keeping the coefficients we have, we need trinomial1 + trinomial2 = 4x -8. notice that the right side does not have a \(\ x^2 \) term which means that the \(\ x^2 \) terms on the left hand side had to add up to zero, a + d = 0. also these sums must be satisfied: b + e = 4 and c + f = -8

OpenStudy (anonymous):

thats it? so i put it in order from the first one right

OpenStudy (anonymous):

@Hero is this right?

hero (hero):

Well, @dpalnc explained it in more general terms. However, you'll need to come up with your own trinomials. For example, if I wanted to create two trinomials that add to get 6x - 4, then I would write the following: \((x^2 + 10x - 6) + (-x^2 - 4x + 2) = 6x - 4\)

hero (hero):

That's just an example

OpenStudy (anonymous):

sooo i cant use his for my question? @Hero

hero (hero):

No, you can't. You have to create your own. Hint: Think outside the box.

OpenStudy (anonymous):

can you just create me one bc i dont understand this and im just trying to get past this @Hero

hero (hero):

Well, if my example didn't help, then I don't know what will

OpenStudy (anonymous):

cant yall just just me what i need not examples? @Hero

hero (hero):

No, we can't